# Higgs Inflation Model with Small Non-Minimal Coupling Constant

### v4

@@ 6.2 Some important properties of normal dynamics
Using eqs.(6.3), (6.4) and (6.14) it is easy to see that $f_{1}(y)\geq 1$. Therefore, the presence of the factor $f_{1}(y)$ in eq.(6.10) leads to an increase in the usual friction force coefficient $3H$. However, due to the factor $(K_{1}-3\tilde{K}_{2}\cdot y)^{-1}\approx(1-1.93\cdot 10^{10}y)^{-1}$ in the last term of eq.(6.10), an analogous effect occurs, leading to a similar strengthening of the usual active force $-U_{eff}^{(tree)\prime}(\varphi)$.
@@
For normal dynamics we can use the upper bound (6.14) on $\dot{\varphi}^{2}$, which together with (6.17) and the condition $\varphi>6M_{P}$ gives an estimate for the third term in eq.(6.10)
$f_{2}\frac{\dot{\varphi}^{2}}{2}\approx 307.8\frac{f_{1}}{M_{P}}\cdot e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}\sqrt{0.17\cdot U_{eff}^{(tree)}}\frac{|\dot{\varphi}|}{\sqrt{2}}<5\cdot 10^{-3}\cdot\frac{f_{1}}{M_{P}}\sqrt{U_{eff}^{(tree)}}\cdot|\dot{\varphi}|.$ (6.18)
@@
$H=\frac{1}{\sqrt{3}M_{P}}\sqrt{K_{eff}+U_{eff}^{(tree)}}>\frac{1}{\sqrt{3}M_{P}}\sqrt{U_{eff}^{(tree)}},$ (6.19)
@@
$|3f_{1}H\dot{\varphi}|>\sqrt{3}\frac{f_{1}}{M_{P}}\sqrt{U_{eff}^{(tree)}}|\dot{\varphi}|.$ (6.20)
Thus, the contribution of $f_{2}\frac{\dot{\varphi}^{2}}{2}$ in eq.(6.10) is at least $2.9\cdot 10^{-3}$ times smaller than $|3f_{1}H\dot{\varphi}|$, and therefore it will be omitted in the further study of normal dynamics.
@@ 6.3 About the slow-roll phase of inflation,
Now, using the results of the previous subsection, we are going to show that the slow-roll approximation is valid during inflation. Furthermore, we want to see what changes occur compared to what we had in section 3.2, where the K-essence structure was completely ignored. Considering the shape of $K_{1}(\varphi)$ and $K_{2}(\varphi)$, one can expect that this effect on inflation at $\varphi>6M_{P}$ is minor. But at the stage when inflation is approaching its end, which happens at $\varphi<6M_{P}$ , where $K_{1}(\varphi)$ and $K_{2}(\varphi)$ noticeably decrease, this influence can be significant.
@@
At the very beginning of inflation, due to the frictional force $\propto 3H\dot{\varphi}$, $\dot{\varphi}^{2}$ decreases, and $\dot{\varphi}^{4}$ decreases even faster. Therefore, not only does $K_{eff}$ remain positive, but the term $\frac{3}{2}\cdot\tilde{K}_{2}\frac{\dot{\varphi}^{4}}{4M_{P}^{4}}$ quickly becomes subdominant compared to $K_{1}\cdot\frac{1}{2}\dot{\varphi}^{2}$. Hence, the first slow-roll condition very quickly reduces to the standard one
@@
Therefore, in the $\varphi$-equation (6.2) one can neglect the contributions $\propto\tilde{K}_{2}(\varphi)\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}$ compared to $K_{1}(\varphi)\approx 1$. Hence, the second slow-roll condition reduces to $|K_{1}(\varphi)\ddot{\varphi}|\ll|3HK_{1}(\varphi)\dot{\varphi}|$, that for $\varphi>6M_{P}$ coincides with the usual one $|\ddot{\varphi}|\ll|3H\dot{\varphi}|$. Thus, the cosmological equations (6.1) and (6.2) in the slow-roll approximation read
$H^{2}=\frac{1}{3M_{P}^{2}}U_{eff}^{(tree)}(\varphi),$ (6.24)
$3HK_{1}(\varphi)\dot{\varphi}=-U_{eff}^{(tree)\prime}(\varphi).$ (6.25)

### v5

@@ 6.2 Some important properties of normal dynamics
Using eqs.(6.3), (6.4) and (6.14) it is easy to see that $f_{1}(y)\geq 1$. Therefore, the presence of the factor $f_{1}(y)$ in eq.(6.10) leads to an increase in the usual friction force coefficient $3H$. However, due to the factor $(K_{1}-3\tilde{K}_{2}\cdot y)^{-1}\approx(1-1.93\cdot 10^{10}y)^{-1}$ in the last term of eq.(6.10), an analogous effect occurs, leading to a similar strengthening of the usual active force $U_{eff}^{(tree)\prime}(\varphi)$.
@@
For normal dynamics we can use the upper bound (6.14) on $\dot{\varphi}^{2}$, which together with (6.17) and the condition $\varphi>6M_{P}$ allow one to obtain an estimate for the third term in eq.(6.10) in a form suitable for comparison with the second term in this equation
$f_{2}\frac{\dot{\varphi}^{2}}{2}\lesssim 307.8\frac{f_{1}}{M_{P}}\cdot e^{-\sqrt{\frac{8}{3}}\frac{\varphi}{M_{P}}}\sqrt{0.17\cdot U_{eff}^{(tree)}}\frac{|\dot{\varphi}|}{\sqrt{2}}<5\cdot 10^{-3}\cdot\frac{f_{1}}{M_{P}}\sqrt{U_{eff}^{(tree)}}\cdot|\dot{\varphi}|.$ (6.18)
@@
$\frac{\dot{a}}{a}=\frac{1}{\sqrt{3}M_{P}}\sqrt{K_{eff}+U_{eff}^{(tree)}}>\frac{1}{\sqrt{3}M_{P}}\sqrt{U_{eff}^{(tree)}},$ (6.19)
@@
$\Big|3f_{1}\frac{\dot{a}}{a}\dot{\varphi}\Big|>\sqrt{3}\frac{f_{1}}{M_{P}}\sqrt{U_{eff}^{(tree)}}|\dot{\varphi}|.$ (6.20)
Thus, the contribution of $f_{2}\frac{\dot{\varphi}^{2}}{2}$ in eq.(6.10) is at least $2.9\cdot 10^{-3}$ times smaller than $|3f_{1}\frac{\dot{a}}{a}\dot{\varphi}|$, and therefore it can be omitted in the further study of normal dynamics. In this approximation equation (6.10) takes the form of the Klein-Gordon equation in the FLRW Universe.
@@ 6.3 About the slow-roll phase of inflation,
Now, using the results of the previous subsection, we are going to show that the slow-roll approximation is valid during inflation. Furthermore, we want to see what changes occur compared to what we had in section 3.2, where the K-essence structure was completely ignored. Considering the shapes of $K_{1}(\varphi)$ and $K_{2}(\varphi)$, one can expect that this effect on inflation at $\varphi>6M_{P}$ is minor. But at the stage when inflation is approaching its end, which happens at $\varphi<6M_{P}$ , where $K_{1}(\varphi)$ and $K_{2}(\varphi)$ noticeably decrease, this influence can be significant.
@@
At the very beginning of inflation, due to the frictional force $\propto 3\frac{\dot{a}}{a}\dot{\varphi}$, $\dot{\varphi}^{2}$ decreases, and $\dot{\varphi}^{4}$ decreases even faster. Therefore, not only does $K_{eff}$ remain positive, but the term $\frac{3}{2}\cdot\tilde{K}_{2}\frac{\dot{\varphi}^{4}}{4M_{P}^{4}}$ quickly becomes subdominant compared to $K_{1}\cdot\frac{1}{2}\dot{\varphi}^{2}$. Hence, the first slow-roll condition very quickly reduces to the standard one
@@
Therefore, in eqs. (6.1) and (6.2) one can neglect the contributions $\propto\tilde{K}_{2}(\varphi)\frac{\dot{\varphi}^{2}}{2M_{P}^{4}}$ compared to $K_{1}(\varphi)\approx 1$. Hence, the second slow-roll condition reduces to $|K_{1}(\varphi)\ddot{\varphi}|\ll|3\frac{\dot{a}}{a}K_{1}(\varphi)\dot{\varphi}|$, that for $\varphi>6M_{P}$ coincides with the usual one $|\ddot{\varphi}|\ll|3\frac{\dot{a}}{a}\dot{\varphi}|$. Thus, the cosmological equations (6.1) and (6.2) in the slow-roll approximation read
$\frac{\dot{a}^{2}}{a^{2}}=\frac{1}{3M_{P}^{2}}U_{eff}^{(tree)}(\varphi),$ (6.24)
$3\frac{\dot{a}}{a}K_{1}(\varphi)\dot{\varphi}=-U_{eff}^{(tree)\prime}(\varphi).$ (6.25)